If a dropped object falls 19.6 m in 2.0 s, how far will it fall in 4.0s?



1
Expert's answer
2021-12-07T20:12:58-0500

The distance is


"d=\frac{gt^2}{2}=\frac{9.8·4^2}{2}=78.4\text{ m}."